OpenStudy (anonymous):

please answer with the solution.. thank you find the radius of a sphere whose surface area and volume are numerically equal?

OpenStudy (anonymous):

So the ratio surface area to volume should be 3/r

OpenStudy (anonymous):

That is the ratio that stands. I'm not really sure about SA=V though. That just doesn't sound right to me

OpenStudy (anonymous):

Surface area is\[4\pi r^2\]Volume is:\[\frac{4}{3}\pi r^3\]Set these equal to each other and solve for r. \[4\pi r^2=\frac{4}{3} \pi r^3\]So,\[r=3\]